A dragster and driver together have mass
888.7 kg . The dragster, starting from rest,
attains a speed of 25.4 m/s in 0.55 s .
Find the average acceleration of the drag-
ster during this time interval.
Answer in units of m/s2
1
Expert's answer
2010-11-08T06:26:10-0500
Fron the equation of the motion
v= v0+at, v0 = 0;
we can find the average acceleration
a = v/t = 25.4/0.55 = 46.2 m/s2